# Double shape quantum phase transitions in the SU3-IBM (I) new $\gamma$-soft phase and the shape phase transition from the new $\gamma$-soft phase to the prolate shape

### v1

@@ IV Conclusion
Based on the existence of the sphere-like spectra Wang22 ; Wang25 ; WangPd , we further discuss the related shape quantum phase transition. In this paper, we have drawn some new conclusions. First, the new $\gamma$-softness is a shape phase, which is very different from previous O(6)-softness as a critical point. Then, we find the double quantum phase transitions along a sing parameter path. We confirm that there is indeed a shape phase transition from the new $\gamma$-soft phase to the prolate shape, and we find experimental evidence that 108Pd is a critical nucleus, which will be studied in detail later. In next paper (II), the scope of the new $\gamma$-soft region in the SU3-IBM for finite-$N$ will be given and the oblate side is also discussed.

### v2

@@ IV Conclusion
Based on the existence of the sphere-like spectra Wang22 ; Wang25 ; WangPd , we further discuss the related shape quantum phase transition. In this paper, we have drawn some new conclusions. First, the new $\gamma$-softness is a shape phase, which is very different from previous O(6)-softness as a critical point. Then, we find the double quantum phase transitions along a sing parameter path. We confirm that there is indeed a shape phase transition from the new $\gamma$-soft phase to the prolate shape, and we find experimental evidence that 108Pd may be a critical nucleus, which will be studied in detail later. In next paper, the scope of the new $\gamma$-soft region in the SU3-IBM for finite-$N$ will be given and the oblate side is also discussed Zhao2 .